Wile E. Coyote is standing at the top of a 300 m cliff. Road Runner startles him, and he falls to the canyon below. Assume that the only acceleration is due to gravity: 9.8 m/sec2 downward.

(a)Write down equations for the position and velocity of the poor coyote.

(b)How much time passes before he hits the ground?

(c)What is his velocity after 4 seconds?
Answer #1

here ,

the height of building , h0 = 300 m

a)

the equation of position ,

from second equation of motion ,

h - h0 = 0 + 0.5 * g * t^2

(h + 300) = 0.5 * 9.8 * t^2

h = 4.9 * t^2 - 300 .....(1)

using first equation of motion

the equation of velocity

v = 0 - g * t

v = - g * t ....(2)

b)

when it hits the ground , h = 0

using (1)

0 = 4.9 * t^2 - 300

t = 7.8 s

c)

at t1 = 4 s

the velocity , v = - g * 4 = - 39.2 m/s

the velocity is 39.2 m/s downward